# Observing network dynamics through sentinel nodes

### v1

@@ Acknowledgements
Appendix S6 Effects of structural features on the inclusion of nodes in optimized node sets
@@
In the main text, we analyzed the degree distributions of optimized sentinel node sets in comparison with the degree distribution of completely randomly node sets. Other structural features of the network may also play roles in making nodes suitable for inclusion in optimized node sets. In this section, we investigate three such features.
S6.1 Effects of three structural features on the approximation error
First, we examined the effect of the average nearest neighbor degrees of the nodes in the optimized node sets. The average nearest neighbor degree of an $i$th node is given by
$k_{{\rm nn},i}=\frac{1}{k_{i}}\sum_{j=1;j\in\mathcal{N}_{i}}^{N}k_{j}$ (S11)
where $\mathcal{N}_{i}$ is the set of the $k_{i}$ neighbors of the $i$th node. We compute the average of $k_{{\rm nn},i}$ over the $n$ nodes in the node set $S$ by
$k_{\rm nn}=\frac{1}{n}\sum_{i=1;i\in S}^{N}k_{{\rm nn},i}$ (S12)
for each optimized and completely random node set. We show in Fig. S11 the distribution of $k_{\rm nn}$ in the optimized and completely random node sets for the coupled double-well dynamics on the ten networks used in Fig. S7. The figure suggests that nodes in optimized node sets tend to have larger, or at least not smaller, $k_{\rm nn}$ than completely random node sets on average. Qualitatively, this observation suggests that our algorithm prefers nodes that “see” more of the network, i.e., nodes with neighbors receiving input from a larger portion of the network.
Figure S11: Distribution of the average nearest neighbor degree of the nodes in sentinel and completely random node sets. The blue and green bars represent optimized and completely random node sets, respectively. We generated 100 node sets of each type to generate each distribution shown. Each panel represents a network. (a) Dolphin. (b) Proximity. (c) Metabolic. (d) Road. (e) Email. (f) ER. (g) BA. (h) HK. (i) GKK. (j) LFR.
Second, we examined the effect of the local clustering coefficient. The local clustering coefficient of an $i$th node is given by [2]
$C_{i}=\frac{\text{number of }i\text{'s neighbors that are adjacent to each other}}{\frac{1}{2}k_{i}(k_{i}-1)}.$ (S13)
We computed the average of $C_{i}$ over the $n$ nodes in node set $S$ by
$C_{\text{local}}=\frac{1}{n}\sum_{i=1;i\in S}^{N}C_{i}$ (S14)
for each optimized and completely random node set. We show the distribution of $C_{\text{local}}$ in the optimized and completely random node sets for the coupled double-well dynamics on various networks in Fig. S12. The figure indicates that $C_{\text{local}}$ does not much differ between optimized and completely random node sets.
Figure S12: Distribution of the average local clustering coefficient in sentinel and completely random node sets. In each panel, we show a binned distribution of $C_{\text{local}}$ for a network. (a) Dolphin. (b) Proximity. (c) Metabolic. (d) Road. (e) Email. (f) ER. (g) BA. (h) HK. (i) GKK. (j) LFR.
@@
Third, we hypothesized that nodes in optimized node sets are more likely to come from different communities than those in completely random node sets. This hypothesis is intuitive because nodes in the same community are expected to show relatively similar dynamics [73, 10, 74, 14], and therefore it is probably redundant to devote multiple sentinel nodes to a single community when one can only use a small number of sentinel nodes. To test the hypothesis, we confined ourselves to six networks, i.e., the five empirical networks and the LFR network, which have relatively strong community structure. We obtained a partition of each network into communities using the Louvain algorithm [75]. Then, for each node set, we counted the number of nodes in the same community and recorded the maximum such number over all the communities, which we denote by $K$. In Fig. S13, we show the distribution of $K$ for the optimized and completely random node sets. We find that the optimized node sets seem to have a somewhat smaller number of sentinel nodes coming from the same community (thus, smaller $K$) than the randomized node sets, supporting our claim, albeit weakly in appearance.
@@
Figure S13: The maximum number of nodes in each node set coming from the same community. (a) Dolphin. (b) Proximity. (c) Metabolic. (d) Road. (e) Email. (f) LFR.
@@
To formalize these observations, we attempted to classify each node set as either optimized or completely random based only on structural features of the node set. A successful classification would imply that one can likely construct a high-performance node set by relying on these node quantities and avoid running our optimization algorithm. We used logistic regression for this classification task. Specifically, we use a generalized linear model with binomial errors and a logit link function in which the dependent variable is the binary outcome of a trial (i.e., either a node set is completely random, which is the reference, or it is optimized). The independent variables are the simulation conditions (i.e., dynamics and network); $k$, the average degree of the $n$ nodes in the node set; $k_{\rm nn}$, the average nearest neighbor degree (i.e., $k_{{\rm nn},i}$) averaged over the $n$ nodes (see Eqs. (S11) and (S12)); $C_{\text{local}}$, the average local clustering coefficient of the $n$ nodes (see Eqs. (S13) and (S14)); and $K$, the maximum number of nodes in a node set assigned to the same community. We analyzed the same number of node sets of each type: 100 optimized node sets for each combination of dynamics and network, and an equal number of completely random node sets (i.e., 400 random node sets for each network). We again exclude the ER network because of its relatively homogeneous degree distribution.
@@
We show the results of this analysis in Table S2. We find that the classification is unsuccessful; our model explains less than 1% of the variance in classification of a node set as random or optimized (McFadden’s pseudo-$R^{2}=0.008$). However, the coefficients on each node quantity are statistically significant and consistent with our analysis above: optimized node sets are associated with smaller $k$ (model coefficient $b=-0.055$, $p=3.62\times 10^{-6}$), larger $k_{\rm nn}$ ($b=0.023$, $p=1.19\times 10^{-4}$), smaller $C_{\text{local}}$ ($b=-1.10$, $p=1.73\times 10^{-4}$), and smaller $K$ ($b=-0.228$, $p<\times 10^{-7}$). Note that the coefficient on $C_{\text{local}}$ is the least extreme among the four ($|z|=3.76$), consistent with our analysis above, while the $p$-value is small due to a large sample size. Additionally, the controls for dynamics in this logistic regression are not significant, suggesting that the composition of optimized node sets may be more strongly related to network structure than dynamics. In sum, our model supports that, across dynamics and networks, optimized node sets avoid large-degree nodes, include nodes with better connected neighbors which do not in turn connect with each other, and avoid nodes that come from the same community. However, all these effects are weak.
@@
Table S2: Regression results in search for features of the optimized node set. The dependent variable is whether the node set is completely random (the reference) or optimized. The model is a logistic regression, i.e., a generalized linear model with binomial errors and a logit link function. SE: standard error, $z$: coefficient estimate divided by the standard error.
@@
       Mutualistic species $0.005$ $0.067$ $0.07$ $0.943$
       SIS $-0.018$ $0.067$ $-0.27$ $0.787$
       Gene-regulatory $0.004$ $0.067$ $0.06$ $0.950$
@@
       Proximity $0.516$ $0.162$ $3.18$ $0.001$
       Metabolic $-0.391$ $0.294$ $-1.33$ $0.183$
       Road $-0.429$ $0.137$ $-3.13$ $0.002$
       Email $0.017$ $0.133$ $0.13$ $0.897$
       BA $-0.529$ $0.141$ $-3.75$ $1.76\times 10^{-4}$
       HK $-0.450$ $0.127$ $-3.54$ $4.07\times 10^{-4}$
       GKK $-0.363$ $0.134$ $-2.72$ $0.007$
       LFR $-0.294$ $0.111$ $-2.65$ $0.008$
$k$ $-0.055$ $0.012$ $-4.63$ $3.62\times 10^{-6}$
$k_{\rm nn}$ $0.023$ $0.006$ $3.85$ $1.19\times 10^{-4}$
$C_{\text{local}}$ $-1.100$ $0.293$ $-3.76$ $1.73\times 10^{-4}$
$K$ $-0.228$ $0.040$ $-5.73$ $<10^{-7}$
Null deviance 9981 $df$: 7199
Residual deviance 9903 $df$: 7184
Pseudo-$R^{2}$ 0.008
S6.2 Heuristic algorithms for node set selection
@@
To explore the possibility of using the information obtained from these analyses to eliminate the need for our optimization algorithm, we designed four heuristic algorithms for node set selection based on our observations. These algorithms do not depend on the dynamics and only use the information on the network structure to different extents. We avoided algorithms using $C_{\text{local}}$ because, as we showed with Fig. S12, its distribution was apparently indifferent between the optimized and completely random node sets for at least some networks.
@@
In the “$k_{\rm nn}$-constrained” algorithm, we reject the bottom 5% of nodes in terms of $k_{\rm nn}$ and select nodes uniformly at random and without replacement from the remaining nodes.
@@
We found that selecting node sets according to these algorithms yielded smaller approximation error than with completely random node sets. Specifically, the $k$-constrained, $k$-quantiled, $k_{\rm nn}$-constrained, and community-based node sets have 1.19, 1.15, 1.30, and 1.17 times smaller approximation error, respectively, than the completely random node sets (see Table S1). This result verifies that the structural properties that our optimization algorithm emphasizes, i.e., degree, average nearest-neighbor degree, and community membership, characterize good sentinel nodes. However, their contribution to suppressing the approximation error is modest. In contrast, one can achieve much smaller approximation error by actually running our optimization algorithm, suggesting that our optimization algorithm determines efficient sentinel node sets by exploiting more complex information on the given network than commonly known properties such as the node degrees and community structure. These observations also hold true when we do not know the test dynamics (Table S3) and for weighted (Tables S6 and S7) and directed (Tables S8 and S9) networks.
@@
We have assumed that we can use the equations of the actual dynamical system in order to optimize node set selection. This condition is unlikely to hold in practice. In this section, we assess the consequences of optimizing with a proposed, or training, dynamics that is not the actual, or test, dynamics. Our procedure works as follows. For a given network, we pretend that we do not know the test dynamics (e.g., SIS dynamics) and therefore run a training dynamics (e.g., coupled double-well dynamics). Then, we analyze the approximation error obtained by the node sets we optimized with the training dynamics when the actual dynamics are the test dynamics (see Figure 5a in the main text).
@@
We followed this procedure for each combination of training dynamics, test dynamics, network, and node set type. Specifically, for each network, we used the 100 node sets optimized on the training dynamics (e.g., coupled double-well dynamics) to calculate approximation errors on each test dynamics (i.e., SIS dynamics, gene-regulatory dynamics, and mutualistic species dynamics if the training dynamics is the coupled double-well dynamics). We carry out this procedure for each pair of different training and test dynamics, each network, and each node set type. Note that only the optimized and degree-preserving node sets use information on the dynamics. Therefore, there are 100 optimized and degree-preserving node sets for each pair of training dynamics and network, but 100 node sets for each network for each of the other five types of node set. As in section S4, we use $\ln\varepsilon$ as the dependent variable and built a generalized linear model with Gaussian errors and an identity link function. We excluded data from the ER network for the same reason as that stated in section S4.

### v3

@@ B Search for node features that would make nodes sentinel
Appendix S10 Approximation error as a function of $n$
In the main text (section IIB), we showed that the degree distributions of optimized node sets remain distinct from those of completely random node sets up to $n=12$ for the dolphin and BA networks. Figure S21 indicates that the approximation errors of optimized node sets also remain distinct from and lower than those of completely random node sets. We note that the approximation error of both optimized and completely random node sets decreases as $n$ increases, which is expected.
Appendix S11 Characterization of sentinel node sets
@@
In the main text, we analyzed the degree distributions of optimized sentinel node sets in comparison with the degree distribution of completely randomly node sets. In section S7, we showed that six features of individual nodes are not predictive of whether or not they are sentinel nodes. However, other features of node sets that are not ascribed to individual nodes’ features may play roles in making node sets suitable as optimized node sets. In this section, we investigate whether or not two such features as well as averages of individual node features over the $n$ nodes in the node set are helpful for constructing good sentinel node sets without resorting to our optimization algorithm.
A Two node set features that cannot be decomposed into single-node features
@@
First, we hypothesized that nodes in optimized node sets are more likely to come from different communities than those in completely random node sets. This hypothesis is intuitive because nodes in the same community are expected to show relatively similar dynamics [105, 10, 106, 14], and therefore it is probably redundant to devote multiple sentinel nodes to a single community when one can only use a small number of sentinel nodes. To test the hypothesis, we confined ourselves to sixteen networks, i.e., the 15 empirical networks and the LFR network, which have relatively strong community structure. We exclude the other four model networks because they do not have community structure by construction. We obtained a partition of each of the 16 networks into communities using the Louvain algorithm [107]. Then, for each node set, we counted the number of nodes in the same community and recorded the maximum such number over all the communities, which we denote by $K$. In Fig. S22, we show the distribution of $K$ for the optimized and completely random node sets. We find that the two node sets are not notably different from each other across the different networks.
@@
Figure S22: The maximum number of nodes in each node set coming from the same community. (a) Dolphin. (b) Proximity. (c) Metabolic. (d) Road. (e) Email. (f) FlyBi. (g) Reactome. (h) Route views. (i) Spanish words. (j) FOLDOC. (k) Tree of life. ($\ell$) English words. (m) Enron. (n) Marker Cafe. (o) Prosper. (p) LFR.
Second, regardless of the community structure, the sentinel node approximation may work better when the members of a sentinel node set are well separated from each other because the different sentinel nodes carry relatively independent information in this case. Therefore, we calculated the distance between two nodes (i.e., the smallest number of hops necessary to reach from one node to the other), which we averaged over all the n(n1)/2n(n-1)/2 node pairs in a node set. We show this average node-to-node distance, denoted by $d$, for all the 20 networks in Fig. S23. We find that sentinel node sets tend to avoid small and large $d$ compared to completely random node sets. Avoidance of small $d$ supports the idea that it may be better for sentinel nodes in the same sentinel node set to be separated from each other. Avoidance of large $d$ may be because sentinel node sets may tend to exclude nodes that are located far from the remainder of the network.
Figure S23: The average distance between node pairs in the sentinel or completely random node set. (a) Dolphin. (b) Proximity. (c) Metabolic. (d) Road. (e) Email. (f) FlyBi. (g) Reactome. (h) Route views. (i) Spanish words. (j) FOLDOC. (k) Tree of life. ($\ell$) English words. (m) Enron. (n) Marker Cafe. (o) Prosper. (p) ER. (q) BA. (r) HK. (s) GKK. (t) LFR.
B Search for node set features that would make node sets sentinel
To formally examine whether or not the $K$ or $d$ value of a node set informs us of the quality of the node set in approximating $\overline{x}$, we first attempted to classify each node set as either optimized or completely random based only on features of the node set. A successful classification would imply that one can likely construct a high-performance node set by relying on these features and avoid running our optimization algorithm. It should be noted that we examined the same question in section S7, but only with features of individual nodes. Because the classification performance was poor when we only use these individual-node features, here we assess the effectiveness of features of node sets. In addition to $K$ and $d$, we used the average of each of the six node features used in section S7 (i.e., $k$, closeness centrality, betweenness centrality, $k_{\text{nn},i}$, $C_{i}$, and coreness) over the $n$ nodes in a node set as additional features of the node set.
@@ B Search for node set features that would make node sets sentinel
We started by using logistic regression for this classification task. Specifically, we use a generalized linear model with binomial errors and a logit link function in which the dependent variable is the binary outcome of a trial (i.e., either a node set is completely random, which is the reference, or it is optimized). The independent variables are the simulation conditions (i.e., dynamics and network); the six node features, each averaged over the $n$ nodes in a node set; $K$, the maximum number of nodes in a node set assigned to the same community; and $d$, the average node-to-node distance in a node set. We analyzed the same number of node sets of each type: 100 optimized node sets for each combination of dynamics and network, and an equal number of completely random node sets (i.e., 400 random node sets for each network). We again exclude the ER network because of its relatively homogeneous degree distribution.
@@
We show the results of this analysis in Table S15. We find that the classification is unsuccessful; our model explains less than 2% of the variance in classification of a node set as random or optimized (McFadden’s pseudo-$R^{2}=0.016$). The $p$ value is small for many independent variables due to a large sample size. Additionally, the controls for dynamics in this logistic regression are not significant, suggesting that the composition of optimized node sets may be more strongly related to network structure than dynamics.
@@
Table S15: Regression results in search for features of the optimized node set. The dependent variable is whether the node set is completely random (the reference) or optimized. The model is a logistic regression, i.e., a generalized linear model with binomial errors and a logit link function. SE: standard error. $z$: coefficient estimate divided by the standard error. LC: local clustering coefficient. The single-node features (i.e., $k$, closeness centrality, betweenness centrality, $k_{\text{nn}}$, LC, and coreness) are averages over the $n$ nodes in the node set. The null deviance is $21,072$ ($df=15,199$) and the residual deviance is $20,726$ ($df=15,170$). The pseudo-$R^{2}$ is $0.016$.
@@
   Mutualistic species $0.011$ $0.046$ $0.242$ $0.809$
   SIS 0.005-0.005 $0.046$ 0.115-0.115 $0.909$
   Gene-regulatory 0.014-0.014 $0.046$ 0.306-0.306 $0.760$
@@
   Proximity $0.269$ $0.139$ $1.941$ $0.052$
   Metabolic 2.452-2.452 $0.231$ 10.625-10.625 $10^{-7}$
   Road $5.974$ $0.455$ $13.131$ $10^{-7}$
   Email $0.130$ $0.140$ $0.928$ $0.353$
   FlyBi $1.558$ $0.189$ $8.246$ $10^{-7}$
   Reactome $1.119$ $0.224$ $4.998$ $5.79\times 10^{-7}$
   Route views $0.144$ $0.148$ $0.975$ $0.329$
   Spanish words 1.556-1.556 $0.202$ 7.691-7.691 $10^{-7}$
   FOLDOC $0.654$ $0.155$ $4.231$ $2.33\times 10^{-5}$
   Tree of life 0.560-0.560 $0.182$ 3.069-3.069 $2.15\times 10^{-3}$
   English words $0.077$ $0.156$ $0.493$ $0.622$
   Enron $0.966$ $0.170$ $5.680$ $10^{-7}$
   Marker Cafe 0.712-0.712 $0.273$ 2.610-2.610 $9.06\times 10^{-3}$
   Prosper $0.074$ $0.218$ $0.340$ $0.734$
   BA $0.819$ $0.163$ $5.026$ $5.00\times 10^{-7}$
   HK $0.844$ $0.159$ $5.300$ $1.16\times 10^{-7}$
   GKK $0.812$ $0.162$ $5.003$ $5.65\times 10^{-7}$
   LFR $3.837$ $0.312$ $12.285$ $10^{-7}$
$k$ 0.025-0.025 $0.003$ 9.659-9.659 $10^{-7}$
Closeness $25.820$ $1.932$ $13.365$ $10^{-7}$
Betweenness 15.740-15.740 $2.531$ 6.220-6.220 $10^{-7}$
$k_{\rm nn}$ $2.04\times 10^{-5}$ $1.27\times 10^{-4}$ 0.161-0.161 $0.872$
LC 0.255-0.255 $0.196$ 1.302-1.302 $0.193$
Coreness $0.030$ $0.005$ $5.569$ $10^{-7}$
$K$ 0.124-0.124 $0.023$ 5.283-5.283 $1.27\times 10^{-7}$
$d$ 0.003-0.003 $0.018$ 0.191-0.191 $0.848$
Although linear regression showed a poor classification result, nonlinear effects of node set features (as opposed to single-node features as investigated in section S7) may make high-quality sentinel node sets. Therefore, we constructed the random forest classifier for the binary classification of node sets using the same data as those used Table S15. We show the classification performance and the variable importance in Table S16. We find a large $R^{2}$ value (i.e., 0.865), indicating that the binary classification is successful. This result is in stark contrast with that with the linear regression shown in Table S15 and implies that nonlinear or combinatorial effects of some independent variables contribute to improving the approximation accuracy of node sets. Table S16 also shows that the average degree over the $n$ nodes, $k$, has the highest variable importance. Therefore, we compare the distribution of $k$ between the optimized and completely random node set in Fig. S24. The figure indicates that sentinel node sets tend to avoid both large $k$ and small $k$. This tendency is a nonlinear effect in terms of $k$, which likely accounts for our result that the random forest, but not the linear regression, performs well at the binary classification. We emphasize that this result on $k$ does not imply that the individual nodes in a good node set are all intermediate-degree nodes. We showed that, while sentinel nodes tend to be non-hubs (see Fig. 3 in the main text and section S7A), different sentinel nodes in an optimized sentinel node set tend to have their degrees scattered (see Fig. 3).
Table S16: Model performance ($R^{2}$) and variable importance for classifying node sets into two categories, completely random or optimized, using a random forest classifier. Close: closeness centrality. Betw: betweenness centrality. LC: local clustering coefficient. Core: coreness. The single-node features (i.e., $k$, closeness centrality, betweenness centrality, $k_{\text{nn}}$, LC, and coreness) are averages over the $n$ nodes in the node set. We used 100 trees and a random subset of three features per node. We used the Gini index as the classification criterion.
$R^{2}$ Dynamics Network $k$ Close Betw $k_{\rm nn}$ LC Core $K$ $d$
$0.865$ $0.031$ $0.020$ $0.420$ $0.232$ $0.142$ $0.140$ $0.052$ $0.349$ $0.004$ $0.065$
Figure S24: The average degree of the nodes in the sentinel or completely random node set. (a) Dolphin. (b) Proximity. (c) Metabolic. (d) Road. (e) Email. (f) FlyBi. (g) Reactome. (h) Route views. (i) Spanish words. (j) FOLDOC. (k) Tree of life. ($\ell$) English words. (m) Enron. (n) Marker Cafe. (o) Prosper. (p) ER. (q) BA. (r) HK. (s) GKK. (t) LFR.
Motivated by this random forest result, we constructed a different random forest model for explaining the approximation error. Because the value of the approximation error in different dynamics are not comparable to each other, we built a random forest model for each dynamics. We show the results in Table S17 for the completely random node sets, merger of the completely random and optimized node sets, and the merge of the completely random, optimized, and degree-preserving random node sets. The results are qualitatively the same as those for the binary classification task shown in Table S16. In other words, in all the three combinations of node sets and all the four dynamics, the random forest model is at least reasonably good at predicting the approximation error of the node set. Furthermore, $k$ remains to be the most important independent variable among the nine independent variables for explaining the approximation error in nine out of twelve random forests shown in the table. Overall, the coreness, closeness centrality, and betweenness centrality are the next most important independent variables after $k$. The last result is also consistent with that for the binary classifier (see Table S16).
Table S17: Model performance ($R^{2}$) and variable importance for random forest models predicting approximation error for different types of node sets. We built separate random forest models for each dynamics, including all networks except the ER network in each random forest model. We used $\ln\varepsilon$ as the dependent variable and squared error as the criterion for finding splits. See the caption of Table S16 for the legends.
Completely random only
Dynamics $R^{2}$ Network $k$ Close Betw $k_{\rm nn}$ LC Core $K$ $d$
Double-well $0.763$ $0.262$ $0.398$ $0.234$ $0.204$ $0.233$ $0.100$ $0.426$ $0.029$ $0.144$
Mutualistic $0.736$ $0.099$ $0.469$ $0.250$ $0.191$ $0.198$ $0.084$ $0.388$ $0.063$ $0.215$
SIS $0.509$ $0.144$ $0.355$ $0.324$ $0.240$ $0.255$ $0.177$ $0.328$ $0.047$ $0.263$
Gene-regulatory $0.731$ $0.117$ $0.805$ $0.115$ $0.265$ $0.308$ $0.063$ $0.184$ $0.019$ $0.057$
Completely random and optimized
Dynamics $R^{2}$ Network $k$ Close Betw $k_{\rm nn}$ LC Core $K$ $d$
Double-well $0.786$ $0.202$ $0.869$ $0.391$ $0.485$ $0.253$ $0.088$ $0.858$ $0.009$ $0.234$
Mutualistic $0.792$ $0.121$ $0.816$ $0.524$ $0.349$ $0.261$ $0.112$ $0.704$ $0.011$ $0.301$
SIS $0.638$ $0.108$ $0.526$ $0.621$ $0.428$ $0.355$ $0.177$ $0.697$ $0.031$ $0.257$
Gene-regulatory $0.917$ $0.086$ $1.118$ $0.181$ $0.444$ $0.209$ $0.075$ $0.349$ $0.010$ $0.083$
Completely random, optimized, and degree-preserving random
Dynamics $R^{2}$ Network $k$ Close Betw $k_{\rm nn}$ LC Core $K$ $d$
Double-well $0.793$ $0.295$ $1.144$ $0.471$ $0.547$ $0.318$ $0.129$ $0.986$ $0.010$ $0.248$
Mutualistic $0.791$ $0.191$ $0.717$ $0.629$ $0.267$ $0.363$ $0.099$ $0.615$ $0.018$ $0.308$
SIS $0.626$ $0.141$ $0.569$ $0.605$ $0.431$ $0.351$ $0.162$ $0.612$ $0.028$ $0.285$
Gene-regulatory $0.829$ $0.142$ $1.478$ $0.280$ $0.574$ $0.228$ $0.107$ $0.547$ $0.015$ $0.098$
C Heuristic algorithms for node set selection
@@
To further explore the possibility of using the information obtained from these analyses to eliminate the need for our optimization algorithm, we investigated four heuristic algorithms for node set selection based on our observations. These algorithms do not depend on the dynamics and only use the information on the network structure to different extents.
@@
In the $k_{\rm nn}$-constrained” algorithm, we reject the bottom 5% of nodes in terms of $k_{\rm nn}$ and select nodes uniformly at random and without replacement from the remaining nodes.
@@
To compare different node selection algorithms, we generated 100 optimized node sets by running the optimization algorithm 100 times, for each combination of dynamics and network. We also generated 100 degree-preserving node sets, 100 completely random node sets for each network, and 100 node sets from each heuristic algorithm (i.e., $k$-constrained, $k$-quantiled, $k_{\rm nn}$-constrained, and community-based) for each network. Note that neither the completely random nor heuristic algorithm node sets depend on the dynamics.
We computed the approximation error for each node set obtained by each of the seven algorithms, each of the four dynamics, and each of the 19 networks, i.e., all but the ER network. We exclude the ER network for the same reason as that for the ANOVA analysis shown in section S4. Then, we carried out a multi-way ANOVA in the same manner as in section S4. The only difference between the two ANOVA analyses is that the current analysis involves seven node sets, whereas the analysis shown in section S4 involves three node set types (i.e., optimized, degree-preserving random, and completely random) that are among the seven node sets.
Overall, our ANOVA model predicts $\ln\varepsilon$ well ($R^{2}=0.72$). In part due to the large sample size (i.e., $53,200$ observations), all three independent variables are highly significant (dynamics: $df=3$, $F=14,278$, $p<10^{-7}$, where the $F$ statistic is the mean-square error of the factor divided by the mean-square error of the residuals; network: $df=18$, $F=245.29$, $p<10^{-7}$; node set type: $df=6$, $F=14,579$, $p<10^{-7}$).
We show the differences between the average error for each pair of node set types, as computed by a Tukey’s honestly significant difference test, in Table S18. As in Table S1, each row of Table S18 specifies the estimated difference in average $\ln\varepsilon$ between the two node set types, the 95% confidence interval of the estimated mean difference, and a $p$ value for the difference adjusted for multiple comparisons. For example, the first row states that, across dynamics and networks, optimized node sets have an average $\ln\varepsilon$ value that is 6.892 smaller than that of completely random node sets. On the natural scale, the average $\varepsilon$ for optimized node sets is $1/e^{-6.892}=984.4$ times smaller than the average $\varepsilon$ for completely random node sets.
Table S18: Differences in average $\ln\varepsilon$ between different node set types when we use the test dynamics for optimization. The differences are computed by a Tukey’s honestly significant difference test. The 95% confidence intervals (CI) of the differences and associated $p$ values, adjusted for multiple comparisons, are also shown.
Difference CI $p$
Optimized - Random 6.892-6.892 [6.983,6.801][-6.983,-6.801] $10^{-7}$
Degree-preserving - Random 3.149-3.149 [3.240,3.058][-3.240,-3.058] $10^{-7}$
$k$-constrained - Random 0.129-0.129 [0.219,0.038][-0.219,-0.038] $5.77\times 10^{-4}$
$k$-quantiled - Random 0.259-0.259 [0.350,0.168][-0.350,-0.168] $10^{-7}$
$k_{\rm nn}$-constrained - Random 0.220-0.220 [0.310,0.129][-0.310,-0.129] $10^{-7}$
Community-based - Random $0.024$ [0.067,0.114][-0.067,0.114] $0.988$
Degree-preserving - Optimized $3.743$ $[3.652,3.834]$ $10^{-7}$
$k$-constrained - Optimized $6.763$ $[6.672,6.854]$ $10^{-7}$
$k$-quantiled - Optimized $6.633$ $[6.542,6.724]$ $10^{-7}$
$k_{\rm nn}$-constrained - Optimized $6.672$ $[6.582,6.763]$ $10^{-7}$
Community-based - Optimized $6.916$ $[6.825,7.006]$ $10^{-7}$
$k$-constrained - Degree-preserving $3.020$ $[2.929,3.111]$ $10^{-7}$
$k$-quantiled - Degree-preserving $2.890$ $[2.799,2.981]$ $10^{-7}$
$k_{\rm nn}$-constrained - Degree-preserving $2.929$ $[2.839,3.020]$ $10^{-7}$
Community-based - Degree-preserving $3.173$ $[3.082,3.263]$ $10^{-7}$
$k$-quantiled - $k$-constrained 0.130-0.130 [0.221,0.039][-0.221,-0.039] $4.68\times 10^{-4}$
$k_{\rm nn}$-constrained - $k$-constrained 0.091-0.091 $182,1.60\times 10^{-4}] 0.0$490.049
Community-based - $k$-constrained $0.152$ $[0.062,0.243]$ $1.52\times 10^{-5}$
$k_{\rm nn}$-constrained - $k$-quantiled $0.039$ [0.051,0.13][-0.051,0.13] $0.863$
Community-based - $k$-quantiled $0.283$ $[0.192,0.373]$ $10^{-7}$
Community-based - $k_{\rm nn}$-constrained $0.243$ $[0.153,0.334]$ $10^{-7}$
@@
The average $\ln\varepsilon$ of any node set type is smaller than that of completely random node sets, except that the community-based algorithm performs poorly, on par with completely random node sets (1st–6th rows of Table S18). In fact, optimized node sets have the smallest error (1st and 7th–11th rows), followed by the degree-preserving node sets (2nd, 7th, and 12th–15th rows), followed by the node sets generated by the heuristic algorithms (remaining rows). The heuristic algorithms, except the community-based algorithm, yielded smaller approximation error than with completely random node sets. Specifically, the $k$-constrained, $k$-quantiled, $k_{\rm nn}$-constrained, and community-based node sets have $1.14$, $1.30$, $1.25$, and $0.98$ times smaller approximation error, respectively, than the completely random node sets (see Table S18). This result verifies that the structural properties that our optimization algorithm emphasizes, i.e., degree, average nearest-neighbor degree, and community membership, characterize good sentinel nodes. However, their contribution to suppressing the approximation error is modest. These observations also hold true for weighted (Table S2) and directed (Table S4) networks.
@@
We have assumed that we can use the equations of the actual dynamical system in order to optimize node set selection. This condition is unlikely to hold in practice. In this section, we assess the consequences of optimizing with a proposed, or training, dynamics that is not the actual, or test, dynamics. Our procedure works as follows. For a given network, we pretend that we do not know the test dynamics (e.g., SIS dynamics) and therefore run a training dynamics (e.g., coupled double-well dynamics). Then, we analyze the approximation error obtained by the node sets we optimized with the training dynamics when the actual dynamics are the test dynamics. See Fig. 5a in the main text. We further show the corresponding results for a weighted mobility-based human network and a directed air-traffic network in Figs. S26 and S27, respectively. The results are similar to those shown in Fig. 5 in the main text, suggesting that our method can work for weighted networks and directed networks and that it is potentially useful for approximating real epidemic outbreaks for which a reasonable dynamics may be unknown.
Figure S25: Transfer learnability of the sentinel node approximation on the Prosper network with $N=89,171$ nodes. (a) Approximation error for optimized node sets evaluated on training and test dynamics. The large markers show the average approximation error over the 100 node sets for each type of node set. (b) Sentinel node approximation for the coupled double-well dynamics on the dolphin network. The states of the $n=\lfloor\ln N\rfloor=11$ sentinel nodes in a sentinel node set are shown in blue. We omit the states of the remaining nodes because there are too many of them. The black line represents the unweighted average state of all nodes, $\overline{x}$. The purple line represents the sentinel node approximation. We show the sentinel node approximation for the (c) mutualistic species, (d) SIS, and (e) gene-regulatory dynamics on the same network when the node set is optimized for the coupled double-well dynamics. The blue lines show $x_{i}^{*}$ for the sentinel nodes used in (b). In (b) and (c), the purple lines cover the black lines.
Figure S26: Transfer learnability of the sentinel node approximation on the weighted version of the Proximity network with $N=410$ nodes. In (b)–(e), we used $n=\lfloor\ln N\rfloor=6$ sentinel nodes. See the caption of Fig. S25 for details.
Figure S27: Transfer learnability of the sentinel node approximation on the Flights network with $N=806$ nodes. In (b)–(e), we used $n=\lfloor\ln N\rfloor=6$ sentinel nodes. See the caption of Fig. S25 for details.
@@
We followed this procedure for each combination of training dynamics, test dynamics, undirected and unweighted network, and node set type. Specifically, for each network, we used the 100 node sets optimized on the training dynamics (e.g., coupled double-well dynamics) to calculate approximation errors on each test dynamics (i.e., SIS dynamics, gene-regulatory dynamics, and mutualistic species dynamics if the training dynamics is the coupled double-well dynamics). We carry out this procedure for each pair of different training and test dynamics, each network, and each node set type. We omit the heuristic algorithms for node set selection considered in section S11C because they do not perform well even when we assume to know the test dynamics (see section S11C for the results). Note that only the optimized and degree-preserving node sets use information on the dynamics. Therefore, there are 100 optimized and degree-preserving node sets for each pair of training dynamics and network, but 100 node sets for each network for completely random node sets.